# Slot Abstractors: Toward Scalable Abstract Visual Reasoning

### v1

@@ 6 Conclusion and Future Directions
In this work, we developed Slot Abstractors, by combining slot-based object-centric encoding mechanisms with Abstractors, a recently proposed approach for implementing the relational bottleneck using a relational variant of cross-attention. We evaluated the Slot Abstractor on four abstract visual reasoning tasks, finding that it achieves overall state-of-the-art systematic generalization of learned abstract rules, and can also be scaled to problems containing a large number of objects and multiple rules.
This work also points to potentially interesting directions for future work. First, although Slot Abstractors demonstrated state-of-the-art performance in many of the OOD generalization regimes of PGM (and achieved state-of-the-art average performance across all regimes), there is still significant room for improvement in some of these regimes. Even for some regimes on which the Slot Abstractor achieved state-of-the-art performance, that performance was sometimes well below what would be expected of human reasoners (e.g., especially for the ‘Extrapolation’ regime). Second, it would be worthwhile to explore how Slot Abstractors would fare in real-world settings (e.g., problems involving real-world images), given recent progress in unsupervised object-centric representation learning for real-world images (Seitzer et al., 2022; Jiang et al., 2023; Wu et al., 2023). Third, the number of slots in slot-based models is fixed, which may pose a challenge for settings in which there are large variations in the number of objects that must be processed. Recent work (Löwe et al., 2023; Stanić et al., 2023) has built methods addressing this limitation by doing away with slots entirely, and future work might explore how such non-slot-based methods can be integrated with relational inductive biases. Finally, future work could also improve the quadratic complexity of Slot Abstractors by using more efficient attention mechanisms (Dao et al., 2022; Dao, 2023).
7 Broader Impact
@@
Appendix A Appendix.
@@
A.1 Datasets
@@ A.1.1 PGM OOD Regimes
A.2 Training Details and Hyperparameters
@@
ART SVRT CLEVR-ART PGM
$H$ 8 8 8 8
$L$ 6 24 24 24
$D_{head}$ 64 64 64 32
$D_{MLP}$ 512 512 512 512

### v2

@@ 6 Conclusion and Future Directions
In this work, we developed Slot Abstractors, by combining slot-based object-centric encoding mechanisms with Abstractors, a recently proposed approach for implementing the relational bottleneck using a relational variant of cross-attention. We evaluated the Slot Abstractor on five abstract visual reasoning tasks including one involving real-world images, finding that it achieves overall state-of-the-art systematic generalization of learned abstract rules, and can also be scaled to problems containing a large number of objects and multiple rules.
Finally, we discuss some limitations of our work and potentially interesting directions for future work. First, although Slot Abstractors demonstrated state-of-the-art performance in many of the OOD generalization regimes of PGM (and achieved state-of-the-art average performance across all regimes), there is still significant room for improvement in some of these regimes. Even for some regimes on which the Slot Abstractor achieved state-of-the-art performance, that performance was sometimes well below what would be expected of human reasoners (e.g., especially for the ‘Extrapolation’ regime). Second, the number of slots in slot-based models is fixed, which may pose a challenge for settings in which there are large variations in the number of objects that must be processed. Recent work (Löwe et al., 2023; Stanić et al., 2023) has built methods addressing this limitation by doing away with slots entirely, and future work might explore how such non-slot-based methods can be integrated with relational inductive biases. Finally, future work could also improve the quadratic complexity of Slot Abstractors by using more efficient attention mechanisms (Dao et al., 2022; Dao, 2023).
Acknowledgements
Shanka Subhra Mondal was supported by Office of Naval Research grant N00014-22-1-2002 during the duration of this work. We would like to thank the reviewers for their valuable feedback, and the Princeton Research Computing, especially William G. Wischer and Josko Plazonic, for their help with scheduling training jobs on the Princeton University Della cluster.
Impact Statement
@@ Acknowledgements
Appendix A Appendix
A.1 Algorithm for Slot Abstractors
@@
A.2 More Experimental Details
For ART and CLEVR-ART tasks, we used a learning rate of $58e-$ 5, batch size of 16, and the number of training epochs are described in Table 15. For SVRT, we used a learning rate of $54e-$ 5, a batch size of 32, and trained for 2000 epochs on each task. We used a single A100 GPU with 80GB memory for training on ART, SVRT, and CLEVR-ART tasks. For training on PGM regimes, we used 4 A100 GPUs with 80GB memory each, a learning rate of $58e-$ 5, and a batch size of 24 per GPU. The number of training epochs for each PGM generalization regime is displayed in Table 16. For training on the V-PROM neutral regime, we also used 4 A100 GPUs with 80GB memory each, a learning rate of $58e-$ 5, and a batch size of 12 per GPU, and trained for 30 epochs. For optimization, we used the ADAM (Kingma & Ba, 2014) optimizer, and for implementation, we used the PyTorch library (Paszke et al., 2017).
A.3 Importance of Slot Attention in DINOSAUR
To justify the importance of the slot attention component, when the DINOSAUR model is used with the Abstractor, we conducted an ablation experiment on the V-PROM neutral regime, by removing the slot attention component from the DINOSAUR model, and used the pre-trained DINO ViT features with the Abstractor. Since, it wasn’t possible to train the ablation model using the patchwise DINO ViT features on an A100 GPU of 80GB memory even with a batch size of 1 (given 9 image panels and 196 patches for each panel, the size of the matrix which models relations between the features would be $1764\times 1764$), and hence we used the CLS output of the DINO ViT features, as feature embeddings $\boldsymbol{z}$ to Abstractor. On V-PROM, the ablation model achieved a test accuracy of 30.01%, compared to 67.7% by our proposed method. This result suggests the importance of slot attention as an effective way of finding a low-dimensional representation, with slots bound to objects in the image.
A.4 Datasets
@@ A.4.1 PGM OOD Regimes
Figure 7: Visual Progressive Matrices (V-PROM) Dataset. Example problems of Raven’s Progressive Matrices involving real-world images. Similar to PGM dataset, the task is to identify the abstract pattern among the images in the first two rows and/or columns of the $3\times 3$ matrix, and use that pattern to fill the missing image in the third row, by selecting one among a set of eight choices.
A.5 Visualization of Slot-wise Reconstruction
Figure 8: Visualization of the individual slot-wise reconstruction for all the 16 slots for an image panel from the PGM dataset.
Figure 9: Visualization of the individual slot-wise reconstruction for all the 16 slots for an image panel from the PGM dataset.
A.6 Training Details and Hyperparameters
@@
Table 10: CNN encoder hyperparameters for V-PROM dataset.
Type Channels Activation Kernel size Stride Padding
Layer Norm - - - - -
1D Conv 768 ReLU 1 1 0
1D Conv 256 - 1 1 0
Table 11: Slot decoder hyperparameters for V-PROM dataset.
Type Channels Activation Kernel size Stride Padding
Spatial Broadcast - - - - -
Position Embedding - - - - -
1D Conv 2048 ReLU 1 1 0
1D Conv 2048 ReLU 1 1 0
1D Conv 2048 ReLU 1 1 0
1D Conv 769 - 1 1 0
@@
Table 13: Slot attention pretraining hyperparameters for V-PROM dataset.
Slot dimensionality $D$ 256
No of iterations $T$ 3
No of slots $K$ 11
Batch size 48
Learning rate $44e-$ 4
Learning rate warmup steps 10000
Epochs 80
@@
ART SVRT CLEVR-ART PGM V-PROM
$H$ 8 8 8 8 8
$L$ 6 24 24 24 24
$D_{head}$ 64 64 64 32 256
$D_{MLP}$ 512 512 512 512 512
Dropout 0 0 0 0 0.1